OpenStudy (anonymous):

Factor Completely: 3x^2+10x-8=

OpenStudy (anonymous):

\[3x^2+10x-8=\]

OpenStudy (anonymous):

(3x-2)(x+4)

OpenStudy (anonymous):

Can you show how you did that?

OpenStudy (anonymous):

i cant because i'm not good with english :)

OpenStudy (anonymous):

it's okay! thanks though :)

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

FOIL- FIRST OUTSIDE INSIDE LAST

OpenStudy (anonymous):

3x^2+10x-8 = y First, one must understand that a polynomial's factors must be of lower term than the polynomial itself. And because of that, any factor for a polynomial must be in the form mx+b one must also notice that if two such factors are multiplied, the resulting polynomial is \[(ax+b)(cx+d) = acx^2+(ad+bc)x+bd\] notice that the polynomial's quadratic coefficient is ac, in our problem, ac is 3. The only factors of 3 are 3 and 1. So, we now have the coefficients of the linear terms of our factors, \[(3x+b)(x+d)\] now, to find b and d, simply find two factors of -8 that will work when you FOIL it out. 8: 1,8; 4,2; Any combination of 1,8 (-1,8; 8,-1; 1,-8; -8,1), when plugged in, will not work to obtain the correct answer. That leaves us with 4,2 to be our answers. If you choose this, \[(3x-2)(x+4)\] ie. you set b = -2, and d = 4, you find that the resulting factors are the correct ones.